# STeP: A General and Scalable Framework for Solving Video Inverse Problems with Spatiotemporal Diffusion Priors

### v1

@@ Appendix B Experimental Details
B.1 Black hole imaging
We introduce the black hole imaging (BHI) problem in more details. In Very Long Baseline Interferometry (VLBI), the cross-correlation of the recorded scalar electric fields at two telescopes, known as the ideal visibility, is related to the ideal source image ${\mathbf{x}}_{0}$ through a 2D Fourier transform, as given by the van Cittert-Zernike theorem [62, 80]. Specifically, the ideal visibility of the $j$-th frame of the target video is
@@ B.1 Black hole imaging
where $(\rho,\delta)$ denotes the angular coordinates of the source image, and $\left(u_{\{a,b\}}^{[j]},v_{\{a,b\}}^{[j]}\right)$ is the dimensionless baseline vector between two telescopes $\{a,b\}$, orthogonal to the source direction.
Due to atmospheric turbulence and instrumental calibration errors, the observed visibility is corrupted by gain error, phase error, and additive Gaussian thermal noise [20, 58]:
@@
where gain errors are denoted by $g_{a}^{[j]},g_{b}^{[j]}$, phase errors are denoted by $\phi_{a}^{[j]},\phi_{b}^{[j]}$, and thermal noise is denoted by ${\mathbf{n}}_{\{a,b\}}^{[j]}$. While the phase of the observed visibility cannot be directly used due to phase errors, the product of three visibilities among any combination of three telescopes, known as the bispectrum, can be computed to retain useful information. Specifically, the phase of the bispectrum, termed the closure phase, effectively cancels out the phase errors in the observed visibilities. Similarly, a strategy can be employed to cancel out amplitude gain errors and extract information from the visibility amplitude [4]. Formally, these quantities are defined as
@@
Here, ($\angle$($\cdot$)∠ ( ⋅ ) denotes the complex angle, and ||$\cdot$|| ⋅ | computes the complex amplitude. For a total of $P$ telescopes, the number of closure phase measurements ${\mathbf{y}}_{\text{cp},\{a,b,c\}}^{[j]}$ at is $\frac{(P-1)(P-2)}{2}$, and the number of log closure amplitude measurements ${\mathbf{y}}_{\text{logca},\{a,b,c,d\}}^{[j]}$ is $\frac{P(P-3)}{2}$, after accounting for redundancy. Since closure quantities are nonlinear transformations of the visibilities, the black hole imaging problem is non-convex.
To aggregate data over different measurement times and telescope combinations, the forward model of black hole imaging for the $j$-th frame can be expressed as
@@
where $\sigma_{\text{cp}}$, $\sigma_{\text{logca}}$, and $\sigma_{\text{flux}}$ are the estimated standard deviations of the measured closure phase, log closure amplitude, and flux, respectively, and $\beta$ is a hyperparameter that controls the strength of the flux regularization, which is empirically determined.
Our BHI experiments are based on the simulation of observing the Sagittarius A black hole with the EHT 2017 array of eight radio telescopes over an observation period of $\approx$100 minutes. We refer the readers to Fig. 5 of [38] for a visualization of the measurement patterns in Fourier space over time. To show the difficulty of this black hole video reconstruction problem, we visualize the dirty images obtained by applying inverse Fourier transform to the ideal visibilities, assuming no measurement errors, in Fig. 7. One can see that substantial spatiotemporal information is lost during the measurement process, so obtaining high-quality reconstructions relies on the effectiveness of incorporating prior information in the reconstruction process.
@@ B.2 Dynamic MRI
MRI is an important imaging technique for clinical diagnosis and biomedical research. Despite its many advantages, MRI is known to be slow because of the physical limitations of the data acquisition in $k$-space. This leads to low patient throughput and sensitivity to patient’s motion [65]. To accelerate the scan speed, instead of fully sampling $k$-space, the compressed subsampling MRI (CS-MRI) technique subsamples $k$-space with masks $\{{\mathbf{m}}^{[j]}\}_{j=1}^{n_{f}}$. In our experiments, the 6×6$\times$6 × acceleration setting with 24 ACS lines leads to $\approx 73\%$ scan time reduction, while the 8×8$\times$8 × acceleration setting with 12 ACS lines leads to $\approx 82\%$ scan time reduction. Fig. 8 visualizes the subsampling masks used in our experiments, where $k_{x},k_{y}$ indicate the frequency encoding and phase encoding directions, respectively. The same mask is applied to the sampling of each individual frame of all videos.
@@ B.3 Baseline Implementations
To make sure we are doing a fair comparison, we implement our two baselines according to Algorithm 1 by modifying several lines. We show the detailed modification below.
IDM+BCS
Follwing [37, 36], we replace the 3D spatiotemporal i.i.d. Gaussian noise in Algorithm 1 to batch consistent Gaussian noise, which is a 3D noise with identical 2D i.i.d. Gaussian frames, as shown in Eq. 17. To implement batch consistency sampling with image diffusion model, we only change the initial nose ${\mathbf{z}}_{t_{N}}$ (line 1 in Algorithm 1) and ${\mathbf{z}}_{t_{i-1}}$ (line 9 in Algorithm 1) from adding 3D spatiotemporal i.i.d. Gaussian noise to batch consistent noise. Moreover, the video diffusion is replaced with an image diffusion model that processes each frame independently.
$\bm{\epsilon}_{\text{BC}}^{[j]}=\bm{\epsilon},\quad\quad\bm{\epsilon}\in\mathbb{R}^{n_{h}\times n_{w}},\forall j=1,2,\cdots,n_{f}$ (17)
IDM
This is by replacing the video diffusion to an image diffusion model that processes each frame independently while keeping the remaining parts changed.
Appendix C Training Details for Video Diffusion Prior
Table 4: Hyper-parameters of the spatiotemporal video diffusion model. We provide and group the hyper-parameters according to each components in the model. The model is trained with 1 NVIDIA A100-SCM4-80GB GPU.
Hyper-parameters Black hole imaging Dynamic MRI
@@ Appendix C Training Details for Video Diffusion Prior
++ image diffusion model training time 5.5h 3.8h
@@
In this section, we show the detail of getting a video diffusion prior on black hole imaging and dynamic MRI, and we summarize the training hyper-parameters in Tab. 4. We define $D_{\text{image}}$ and $D_{\text{video}}$ as the image and video datasets, containing $N_{\text{image}}$ and $N_{\text{video}}$ data points, respectively. The image dataset $D_{\text{image}}$ includes all individual frames from the video dataset $D_{\text{video}}$, along with additional large-scale image data to enhance generalization. For data augmentation, we apply random horizontal/vertical flipping and random zoom-in-and-out to improve robustness and diversity in training.
We first train the compression functions, the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, on an image dataset. The training objective consists of an L1 reconstruction loss combined with a KL divergence term scaled by a factor $\beta_{KL}$. The loss function for training is as defined in Eq. 18. The Adam optimizer is used as the default optimizer throughout the paper. The loss function for training the variational autoencoder (VAE) is given by:
@@
where $p({\mathbf{z}}_{0})$ is the standard Gaussian $\mathcal{N}({\bm{0}},{\bm{I}})$ and $q_{\phi}({\mathbf{z}}_{0}|{\mathbf{x}}_{0})$ is the isotropic Gaussian distribution over ${\mathbf{z}}_{0}$ where the mean and standard deviation is given by $\mathcal{E}({\mathbf{x}}_{0})$. Next, we train the image diffusion UNet ${\mathbf{s}}_{{\bm{\theta}}}$ using the standard score-matching loss, as defined in Eq. 19, following [27, 54].
@@
After pertaining, the image diffusion UNet ${\mathbf{s}}_{{\bm{\theta}}}$ is then converted to a spatiotemporal UNet by adding zero-initialized temporal modules to 2D spatial modules and fine-tune jointly with video and image datasets. We use the same Eq. 19 without change ${\mathbf{x}}_{0}$ to video or pseudo video input and use the encoder to process each frame independently.
After pretraining, the image diffusion U-Net ${\mathbf{s}}_{{\bm{\theta}}}$ is transformed into a spatiotemporal UNet by integrating zero-initialized temporal modules into the existing 2D spatial modules. The model is then fine-tuned jointly using both video and image datasets. We use the same loss as in Eq. 19, by changing ${\mathbf{x}}_{0}$ to a video or a pseudo-video input. Each frame is independently processed using the encoder $\mathcal{E}$, ensuring that spatial representations remain aligned while temporal consistency is learned through the added temporal modules.
Appendix D Discussion
D.1 Sampling Efficiency
We discuss the sample efficiency in this section. The sampling time of STeP  depends on the total number of video diffusion model calling $N_{\text{vdm}}$ and total number of decoder, and its gradient calling $N_{\text{dec}}$. We summarize these parameters and sampling requirement in Tab. 5.
Table 5: The sampling requirement and number of function callings in STeP for two problems. The run time and memory is tested using 1 NVIDIA A 100-SCM4-80GB GPU.
$N_{\text{dec}}$ $N_{\text{vdm}}$ time (s) memory (GB)
Black hole imaging 1500 500 645 52
Dynamic MRI 1060 400 332 23
D.2 Limitations and Future Extension
Though STeP is a general and scalable framework for solving video inverse problem with spatiotemporal diffusion prior, the sampling cost of STeP is high due to the requirement of backpropagation through decoder $\mathcal{D}$ in MCMC updates in Algorithm 1. This is forcing us to balance between the capability of the decoder and its computational cost. We leave the exploration of performing MCMC updates in pixel space or other approaches to bypass calling decoder as future work.
Appendix E More Results & Visualization

### v2

@@ Appendix B Experimental details
B.1 Black hole video reconstruction
We introduce the black hole video reconstruction problem in more detail. The goal is to reconstruct a video ${\mathbf{x}}_{0}\in\mathbb{R}^{n_{f}\times n_{h}\times n_{w}}$ of a rapidly moving black hole. Each measurement, or visibility, is given by correlating the measurements from a pair of telescopes to sample a particular spatial Fourier frequency of the source with very long baseline interferometry (VLBI) (van1934wahrscheinliche, ; zernike1938concept, ). In VLBI, the cross-correlation of the recorded scalar electric fields at two telescopes, known as the ideal visibility, is related to the ideal source video ${\mathbf{x}}_{0}$ through a 2D Fourier transform, as given by the van Cittert-Zernike theorem (van1934wahrscheinliche, ; zernike1938concept, ). Specifically, the ideal visibility of the $j$-th frame of the target video is
@@ B.1 Black hole video reconstruction
where $(\rho,\delta)$ denotes the angular coordinates of the source video frame, and $\left(u_{\{a,b\}}^{[j]},v_{\{a,b\}}^{[j]}\right)$ is the dimensionless baseline vector between two telescopes $\{a,b\}$, orthogonal to the source direction.
Due to atmospheric turbulence and instrumental calibration errors, the observed visibility is corrupted by gain error, phase error, and additive Gaussian thermal noise (m87paperiii, ; sun2024provable, ):
@@
where gain errors are denoted by $g_{a}^{[j]},g_{b}^{[j]}$, phase errors are denoted by $\phi_{a}^{[j]},\phi_{b}^{[j]}$, and thermal noise is denoted by ${\mathbf{n}}_{\{a,b\}}^{[j]}$. While the phase of the observed visibility cannot be directly used due to phase errors, the product of three visibilities among any combination of three telescopes, known as the bispectrum, can be computed to retain useful information. Specifically, the phase of the bispectrum, termed the closure phase, effectively cancels out the phase errors in the observed visibilities. Similarly, a strategy can be employed to cancel out amplitude gain errors and extract information from the visibility amplitude (Blackburn_2020, ). Formally, these quantities are defined as
@@
Here, ($\angle$($\cdot$)∠ ( ⋅ ) denotes the complex angle, and ||$\cdot$|| ⋅ | computes the complex amplitude. For a total of $P$ telescopes, the number of closure phase measurements ${\mathbf{y}}_{\text{cp},\{a,b,c\}}^{[j]}$ at is $\frac{(P-1)(P-2)}{2}$, and the number of log closure amplitude measurements ${\mathbf{y}}_{\text{logca},\{a,b,c,d\}}^{[j]}$ is $\frac{P(P-3)}{2}$, after accounting for redundancy. Let $D^{[j]}_{\text{cp}}$ and $D^{[j]}_{\text{camp}}$ to indicate the dimension of ${\mathbf{y}}_{\text{cp}}^{[j]}$ and ${\mathbf{y}}_{\text{logca}}^{[j]}$. Since closure quantities are nonlinear transformations of the visibilities, the black hole video reconstruction problem is non-convex.
To aggregate data over different measurement times and telescope combinations, the forward model of black hole video reconstruction for the $j$-th frame can be expressed as
@@
$-\log p({\mathbf{y}}\mid{\mathbf{x}}_{0})$ $\propto\underbrace{\sum_{j=1}^{n_{f}}\dfrac{1}{n_{f}D_{\text{cp}}^{[j]}\sigma_{\text{cp}}^{2}}\left\|\mathcal{A}^{[j]}_{\text{cp}}({\mathbf{x}}_{0})-{\mathbf{y}}^{[j]}_{\text{cp}}\right\|^{2}}_{\bm{\chi}_{\text{cp}}^{2}}+\underbrace{\sum_{j=1}^{n_{f}}\dfrac{1}{n_{f}D_{\text{logca}}^{[j]}\sigma_{\text{logca}}^{2}}\left\|\mathcal{A}^{[j]}_{\text{logca}}({\mathbf{x}}_{0})-{\mathbf{y}}^{[j]}_{\text{logca}}\right\|^{2}}_{\bm{\chi}_{\text{logca}}^{2}}$
$+\beta\underbrace{\sum_{j=1}^{n_{f}}\dfrac{1}{n_{f}\sigma_{\text{flux}}^{2}}\left\|\mathcal{A}^{[j]}_{\text{flux}}({\mathbf{x}}_{0})-{\mathbf{y}}^{[j]}_{\text{flux}}\right\|^{2}}_{\bm{\chi}_{\text{flux}}^{2}}$ (12)
where $\sigma_{\text{cp}}$, $\sigma_{\text{logca}}$, and $\sigma_{\text{flux}}$ are the estimated standard deviations of the measured closure phase, log closure amplitude, and flux, respectively, and $\beta$ is a hyperparameter that controls the strength of the flux regularization, which is empirically determined. To evaluate the data fitting, we introduce a unified $\bm{\tilde{\chi}}^{2}$ statistics
The $\tilde{\bm{\chi}}^{2}$ is no less than $1$ and closer to $1$, indicating better measurement data fit. In our experiments we use an average $(\tilde{\bm{\chi}}^{2}_{\text{cp}}+\tilde{\bm{\chi}}^{2}_{\text{camp}})/2$ for the data misfit metrics for evaluation.
Our experiments are based on the simulation of observing the Sagittarius A black hole with the EHT 2017 array of eight radio telescopes over an observation period of around 100 minutes. We refer the readers to Fig. 5 of (levis2021fluid, ) for a visualization of the measurement patterns in Fourier space over time. To show the difficulty of this black hole video reconstruction problem, we visualize the dirty video frames obtained by applying the inverse Fourier transform to the ideal visibilities, assuming no measurement errors, in Figure 8. One can see that substantial spatiotemporal information is lost during the measurement process, so obtaining high-quality reconstructions relies on the effectiveness of incorporating prior information in the reconstruction process.
@@ B.2 Dynamic MRI
MRI is an important imaging technique for clinical diagnosis and biomedical research, where the objective is to recover a video ${\mathbf{x}}_{0}\in\mathbb{C}^{n_{f}\times n_{h}\times n_{w}}$ of the heart from the subsampled Fourier space (a.k.a $k$-space) measurements ${\mathbf{y}}$. Despite its many advantages, MRI is known to be slow because of the physical limitations of the data acquisition in $k$-space. This leads to low patient throughput and sensitivity to patient’s motion (wang2024cmrxrecon, ). To accelerate the scan speed, instead of fully sampling $k$-space, the compressed subsampling MRI (CS-MRI) technique subsamples $k$-space with masks $\{{\mathbf{m}}^{[j]}\}_{j=1}^{n_{f}}$. Mathematically, this can be formulated as
${\mathbf{y}}^{[j]}={\mathbf{m}}^{[j]}\odot\mathcal{F}\left({\mathbf{x}}_{0}^{[j]}\right)+{\mathbf{n}}^{[j]}\in\mathbb{C}^{n}\quad\text{for}j=1,...,n_{f},$ (14)
where ${\mathbf{m}}^{[j]}\in\{0,1\}^{n_{h}\times n_{w}}$ is the subsampling mask for the $j$-th frame, $\odot$ denotes element-wise multiplication, $\mathcal{F}$ is the Fourier transform, and ${\mathbf{n}}^{[j]}$ is the measurement noise. In our experiments, we used subsampling masks with an equi-spaced pattern (similar to those visualized in wang2024cmrxrecon ) of both 6×6$\times$6 × acceleration with 24 auto-calibration signal (ACS) lines (Tab. 4) and 8×8$\times$8 × acceleration with 12 ACS lines (Tab. 1). For dynamic MRI, we use the Gaussian likelihood function
$-\log p({\mathbf{y}}|{\mathbf{x}}_{0})\propto\|\mathcal{A}({\mathbf{x}}_{0})-{\mathbf{y}}\|^{2}_{2}.$ (15)
Fig.9 visualizes the subsampling masks used in our experiments, where $k_{x},k_{y}$ indicate the frequency encoding and phase encoding directions, respectively. The same mask is applied to the sampling of each frame of all videos.
@@ B.3 Baseline implementations
In this section, we explain the detailed implementation of our baseline methods. We follow the same grouping in Sec. 4.1.
Group 1: simple heuristics
We implement batch independent sampling (BIS) and batch consistent sampling (BCS) following (kwon2025solving, ; kwon2024visionxlhighdefinitionvideo, ). BIS is implemented to reconstruct each frame independently, whereas BCS promotes temporal consistency by using identical ("batch-consistent") noise across the temporal dimension. To ensure a fair comparison, we adapt Algorithm 1 by initializing the noise variables ${\mathbf{z}}_{t_{N}}$ (line 1) and ${\mathbf{z}}_{t_{i-1}}$ (line 9) with batch-consistent noise.
Group 2: noise warping
We follow (deng2025infiniteresolution, ; chang2024how, ) in performing noise warping based on optical flow estimated from measurements. As described in Sec. 4.1, we derive optical flow using ground-truth black hole videos and inverse Fourier-transformed dynamic MRI videos with a pretrained model (teed2020raftrecurrentallpairsfield, ). Since we utilize a latent diffusion model, we further downsample the optical flow via interpolation to match the latent noise dimension. We adapt Algorithm 1 accordingly by replacing the original i.i.d. noise with the warped noise while keeping other components unchanged. We implement Bilinear, Bicubic, and Nearest warping strategies using their corresponding interpolation methods, and follow the implementation from https://github.com/yitongdeng-projects/infinite_resolution_integral_noise_warping_code for $\int$-noise. Lastly, since (daras2024warped, ) does not provide publicly available code, we implement GP-Warp following Equation (2) from their paper.
Table 3: Summary of the training of spatiotemporal diffusion prior. We provide and group the hyperparameters according to each component in the model. The model is trained with 1 NVIDIA A100-SCM4-80GB GPU.
Hyper-parameters Black hole video reconstruction Dynamic MRI
@@
image diffusion model training time 5.5h 3.8h
@@
Appendix C Training details for spatiotemporal diffusion prior
In this section, we show the detail of getting a video diffusion prior on black hole video reconstruction and dynamic MRI, and we summarize the training hyper-parameters in Table 3. We define $D_{\text{image}}$ and $D_{\text{video}}$ as the image and video datasets, containing $N_{\text{image}}$ and $N_{\text{video}}$ data points, respectively. The image dataset $D_{\text{image}}$ includes all individual frames from the video dataset $D_{\text{video}}$, along with additional large-scale image data to enhance generalization. For data augmentation, we apply random horizontal/vertical flipping and random zoom-in-and-out to improve robustness and diversity in training.
We first train the compression functions, the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, on an image dataset. The training objective consists of an L1 reconstruction loss combined with a KL divergence term scaled by a factor $\beta_{KL}$. The loss function for training is as defined in Equation (16). The Adam optimizer is used as the default optimizer throughout the paper. The loss function for training the variational autoencoder (VAE) is given by:
@@ Appendix C Training details for spatiotemporal diffusion prior
where $p({\mathbf{z}}_{0})$ is the standard Gaussian $\mathcal{N}({\bm{0}},{\bm{I}})$ and $q_{\phi}({\mathbf{z}}_{0}|{\mathbf{x}}_{0})$ is the isotropic Gaussian distribution over ${\mathbf{z}}_{0}$ where the mean and standard deviation is given by $\mathcal{E}({\mathbf{x}}_{0})$. Next, we train the image diffusion UNet ${\mathbf{s}}_{{\bm{\theta}}}$ using the standard score-matching loss, as defined in Equation (17), following (ho2020denoising, ; song2021scorebased, ).
@@
After pertaining, the image diffusion UNet ${\mathbf{s}}_{{\bm{\theta}}}$ is then converted to a spatiotemporal UNet by adding zero-initialized temporal modules to 2D spatial modules and fine-tune jointly with video and image datasets. We use the same Equation (17) without change ${\mathbf{x}}_{0}$ to video or pseudo video input and use the encoder to process each frame independently.
After pretraining, the image diffusion U-Net ${\mathbf{s}}_{{\bm{\theta}}}$ is transformed into a spatiotemporal UNet by integrating zero-initialized temporal modules into the existing 2D spatial modules. The model is then fine-tuned jointly using both video and image datasets. We use the same loss as in Equation (17), by changing ${\mathbf{x}}_{0}$ to a video or a pseudo-video input. Each frame is independently processed using the encoder $\mathcal{E}$, ensuring that spatial representations remain aligned while temporal consistency is learned through the added temporal modules.
Appendix D Limitations
Though STeP is a general framework for solving scientific VIPs with spatiotemporal diffusion prior, the sampling cost of STeP is relatively high due to the requirement of backpropagation through decoder $\mathcal{D}$ in MCMC updates in Algorithm 1. So we have to strike a balance between the capability of the decoder and its computational cost. We leave the exploration of performing MCMC updates in pixel space or other better approaches to bypass backpropagation through the decoder as future work.
Appendix E Broader impacts
Our work introduces a new method for solving scientific video inverse problems, with potential applications in astronomy, medical imaging, and physics, where accurate reconstruction from sparse data is crucial. However, the learned spatiotemporal prior may introduce biases; thus, practitioners should responsibly evaluate the reconstructions and mitigate any potential negative impacts.
Appendix F Licenses
We list the licenses of all the assets we used in this paper:
Appendix G More results & visualization
Table 4: Additional results on dynamic MRI with 6$\times$ acceleration. Following the same setup as in Tab. 1, we report the quantitative results on 10 test videos.
Tasks Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) Data Misfit ($\downarrow$)
MRI (6$\times$) BIS kwon2025solving 39.47 (0.59) 0.958 (0.007) 0.086 (0.011) 43.26 (1.23) 0.962 (0.005) 113.17 11.071 (0.740)
BCS kwon2025solving 40.69 (0.57) 0.959 (0.006) 0.081 (0.012) 44.73 (1.31) 0.974 (0.004) 110.01 11.085 (0.720)
Bilinear chang2024how 40.85 (0.57) 0.960 (0.006) 0.080 (0.012) 44.84 (1.28) 0.975 (0.004) 114.37 11.038 (0.735)
Bicubic chang2024how 40.71 (0.67) 0.959 (0.007) 0.079 (0.012) 44.74 (1.38) 0.974 (0.005) 106.82 11.068 (0.755)
Nearest chang2024how 40.37 (0.56) 0.960 (0.007) 0.080 (0.012) 44.81 (1.41) 0.974 (0.005) 110.91 11.050 (0.739)
$\int$-noise chang2024how ; deng2025infiniteresolution 40.09 (0.50) 0.960 (0.006) 0.082 (0.012) 44.77 (1.34) 0.974 (0.005) 111.92 11.059 (0.731)
GP-Warp daras2024warped 39.50 (0.48) 0.959 (0.007) 0.080 (0.012) 44.53 (1.33) 0.973 (0.005) 105.70 11.070 (0.727)
STeP (video only) 40.76 (0.43) 0.967 (0.005) 0.077 (0.012) 46.38 (1.82) 0.981 (0.005) 101.83 10.788 (0.713)
STeP (image-video joint) 41.39 (0.52) 0.969 (0.005) 0.076 (0.012) 46.61 (1.72) 0.982 (0.004) 98.15 10.808 (0.723)
Data misfit values for samples in Fig. 1.
We report the data misfit values in Tab. 5.
Table 5: The data misfit values for samples shown in Fig. 1.We report the data misfit metrics for the three obtained modes, which demonstrate that all modes fit the measurement data equally well.
Metrics Mode 1 Mode 2 Mode 3
$\bm{\chi}^{2}_{\text{cp}}$ 1.045 0.987 1.084
$\bm{\chi}^{2}_{\text{logca}}$ 1.007 1.001 1.202
Data Misfit 1.026 1.007 1.143